OpenStudy (anonymous):

When using the substitution method, what expression for y in the first equation would be substituted in the second equation? 5x – y = –2 9x + 2y = 23

OpenStudy (radar):

y=5x+2

OpenStudy (anonymous):

thank you!

OpenStudy (anonymous):

Which of the below systems of equations matches the following graph? Answer x + 3y = 9 2x – 3y = 0 x – 3y = 9 2x + 3y = 0 x – 3y = –9 2x – 3y = 0 x + 3y = –9 2x + 3y = 0